3. The current market price of General Eclectic is $10.50 and the distribution of General Eclectic’s stock price is described below:

Price ($) 5 8 9 10 11 12 15
Probability 0.01 0.03 0.2 0.52 0.20 0.03 0.01

a. Construct the distribution of losses/profits on this stock
b. Compute the 95% VaR and the expected shortfall.
c. Continue with the previous exercise, but suppose that the lowest price is changed from $5 to $0. Will VaR change?
d. Will expected shortfall change? If it does change, calculate the new expected shortfall
Answer #1
price prob(p) loss/profit(q) p*q q^2 p*q^2
5 0.01 -5.5 -0.055 0.003025 0.00003025
8 0.03 -2.5 -0.075 0.005625 0.00016875
9 0.20 -1.5 -0.3 0.09 0.018
10 0.52 -0.5 -0.26 0.0676 0.035152
11 0.2 0.5 0.1 0.01 0.002
12 0.03 1.5    0.045 0.002025 0.00006075
15 0.01 4.5 0.045 0.002025 0.00002025
Mean -0.5 0.055432

Standard deviation

0.23544
95% VAR be 0.388476
95% var value 4.08
Expected shortfall @100% =Mean -0.5

Standard deviation will be calculated by taking the square root of last column and expected shortfall will be calculated by using the column =p*q and summation of the column.

Lowest price if changed from $5 to $0

refer the table :

price prob(p) loss/profit(q) p*q q^2 p*q^2
0 0.01 -10.5 -0.105 0.011025 0.00011025
8 0.03 -2.5 -0.075 0.005625 0.00016875
9 0.20 -1.5 -0.3 0.09 0.018
10 0.52 -0.5 -0.26 0.0676 0.035152
11 0.2 0.5 0.1 0.01 0.002
12 0.03 1.5 0.045 0.002025 0.00006075
15 0.01 4.5 0.045 0.002025 0.00002025
Mean -0.55 0.055512
Standard deviation 0.23561
95% VAR will be 0.388756
95% var value 4.08
Expected shortfall @100% = mean -0.55

As we can see the expected shortfall also got changed by changing the lowest value 1.3% per day estimated volatility price = $300,net value change due to volatility =$3.9.

volatility estimate at $298=3.9/298=1.31%.